OpenStudy (anonymous):

Y=x^2-6x+1 x=3,-8?? Am I right?

OpenStudy (anonymous):

\[\frac{ -b \sqrt{b ^{2}-4ac}}{ 2a }\] a=1 b=-6 c=1 Therefore: \[\frac{ -(-6)\sqrt{(-6)^{2}-4(1)(1)} }{ 2(1) }\] So: \[\frac{ 6\sqrt{36-4} }{ 2 }\] \[\frac{ 6+\sqrt{32} }{ 2}\] and \[\frac{ 6-\sqrt{32} }{ 2 }\] \[\sqrt{32} \approx 5.657\]

OpenStudy (anonymous):

\[6-\sqrt{32} \approx .343\] and \[6+5.657 \approx 11.657\] Divide both answers by 2. \[x \approx 5.829\] or \[x \approx .1715\] Keeping the radical gives an exact answer.

OpenStudy (anonymous):

Oh wow, so I was way off.

OpenStudy (anonymous):

yeah it says use radicals as needed.

OpenStudy (anonymous):

It happens! I hope the way I laid out the steps was clear enough!